The cross-sectional area of a wire containing 0.67 A is 4.9 × 10^–7 m^2 . If there are 8.5 × 10^28 charge carriers per m^3 in this wire, what is the drift speed of the charge carriers?  


1
Expert's answer
2021-12-20T10:27:22-0500

We can find the drift speed of the charge carriers as follows:


"v_d=\dfrac{I}{nAq},""v_d=\dfrac{0.67\ A}{8.5\times10^{28}\ \dfrac{el}{m^3}\times4.9\times10^{-7}\ m^2\times1.6\times10^{-19}\ C},""v_d=1.0\times10^{-4}\ \dfrac{m}{s}."